OpenStudy (anonymous):

can someone help me with a question. i did all the work i just need to know if my answer is correct

OpenStudy (imstuck):

What do you need?

OpenStudy (anonymous):

this is my question Jordan is a manager of a car dealership. He has 3 professional car washers to clean the entire lot of cars, Jennifer, Arianna, and Matthew. Jennifer can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Matthew is new to the car dealership, so no one knows how long it will take him. Jordan assigns all of them to wash the cars together. Explain to Jordan how this task can tell him how long it would take Matthew to complete the task if he worked by himself. Use complete sentences.

OpenStudy (anonymous):

the last line that i wrought down is M = 154/154- 25T

OpenStudy (imstuck):

Ok, this is going to be kinda tough to write it out and explain it properly but I will try.

OpenStudy (anonymous):

i know what you mean, this is a long equation. is what i wrote for the last line good?

OpenStudy (imstuck):

If all of them work together and the manager finds that it takes them 8 hours to get all the cars washed, for example, the manager would use this equation to find out how long it would take matt to wash the cars alone.\[Jennifer + Arianna + Matt = 8 hours\]or\[\frac{ 1 }{ 14 }+\frac{ 1 }{11 }+\frac{ 1 }{ t }=\frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

how did you get 8?

OpenStudy (imstuck):

Going from the first equation to the second you are figuring how much work can be done by each per hour. Hence the reason for the 1/ fractions. Solve this for t and that will tell you how much work Matt can get done in 1 hour. Then readjust the fraction to give you the whole job's number of hours.

OpenStudy (imstuck):

8 was just an example to use.

OpenStudy (imstuck):

So if you were to explain it in words, it would be like this, ok?

OpenStudy (imstuck):

Have the 3 workers do the washing together and log the total number of hours it takes all three of them to complete the job. Then add 1/14 (the work Jennifer can get done in 1 hour) to 1/11 (the work Arianna can get done in 1 hour) to 1/t ( the mystery number of how much work Matt can get done in 1 hour) and set that equal to 1/the total number of hours of work it took all of them to do the job. Find the LCD among the denominators and solve the resulting equation for t, the amount of work Matt can get done in hour.

OpenStudy (imstuck):

That's how you would do it. If you had a total for the number of hours it took all 3 of them, we could figure out Matt's time.

OpenStudy (anonymous):

ok well the LCD =154

OpenStudy (imstuck):

Where did you get 154?

OpenStudy (anonymous):

11 goes into 154 14 times

OpenStudy (imstuck):

That can't be the LCD; the LCD will contain a variable.

OpenStudy (imstuck):

It will contain whatever variable you designate as Matt's number of work hours.

OpenStudy (anonymous):

so do you think this is worng https://answers.yahoo.com/question/index?qid=20140105181818AAkkZrL

OpenStudy (imstuck):

I see the logic, but it's not a definitive answer. It still doesn't give you a definite number for Matt's hours. You can't find it because you don't know how long it takes them to work all together. It's not wrong; that's the work all worked out. You said you needed sentences to explain how to find Matt's time. That's what I did for you. If you notice I did mention to find the common denominator among all the denominators. That's where this person got the 154, by multiplying 11 and 14. But your question wanted a written explanation not a mathematical one.